OpenStudy (anonymous):

The legs of a right triangle are 3 units and 5 units. What is the length of the hypotenuse? Round your answer to the nearest hundredth. 4.00 units 2.83 units 5.83 units 8.00 units I'm pretty sure its 8 but i just need conformation. I just added up both sides but idk if thats correct because hypotenuse is the longest side so it could possibly 5

OpenStudy (jdoe0001):

 \(\bf \textit{pythagorean theorem}=c^2=a^2+b^2\to c=\pm\sqrt{a^2+b^2}\)

OpenStudy (anonymous):

thank you, i got 8.3 and rounded its 8 haha